# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

### v4

@@ A.4.4 Hyperparameters for Instruction-Following RL Agents
- chans [64, 128, 128]
- outsize 256
- nblock 2
- post_pool_groups 1
- init_norm_kwargs
- batch_norm false
- group_norm_groups 1
@@
- layer_norm true
Table 5: Crafter and Minigrid RL Parameters
@@
- ppo_nepoch 3
- ppo_nbatch 8
- clip_param 0.2
- vf_loss_coef 0.5
- ent_coef 0.01
- lr 3.0e-4
- max_grad_norm 0.5
- aux_freq 8
- aux_nepoch 6
- pi_dist_coef 1.0
- vf_dist_coef 1.0
Table 6: Montezuma RL Training Parameters
@@
- update_proportion 0.25
- ppo_nepoch 3
- ppo_batch_size 256
- clip_param 0.1
- vf_loss_coef 0.5
- ent_coef 0.001
- lr 1.0e-4
A.5 Additional Details of the Experiments on the Impact of Noisy Rewards
Evaluation Metric Details In our experiments, we used a score metric adapted from the Crafter paper to evaluate agent performance across different environments. This score metric aggregates success rates for individual subtasks using a geometric mean. Formally, the score metric is defined as follows:
  1. 1.
  2. 2.
  3. 3.
A.5.1 Details of Manipulated Trajectory-Instruction Pairs to Evaluate Robustness
We evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. These manipulations were designed to test the robustness of the models against various types of noise. Here’s a detailed breakdown of the manipulations:
  1. 1.
  2. 2.
  3. 3.
    Instruction Rephrasing: We rephrase the original instructions while maintaining their core meaning. This evaluates the model’s robustness to linguistic variations and its ability to capture the essential semantic content of instructions.
  4. 4.
    Concatenation and Order Swapping: Given two trajectory-instruction pairs $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we created concatenated pairs and then swapped the order in one modality. For example:
    • Original concatenation: $(\tau_{1}+\tau_{2},l_{1}+l_{2})$
    • Swapped trajectory: $(\tau_{2}+\tau_{1},l_{1}+l_{2})$
    • Swapped instruction: $(\tau_{1}+\tau_{2},l_{2}+l_{1})$
    This tests the model’s sensitivity to the order of components in multi-step tasks.
  5. 5.
    Concatenation with Partial Content: We concatenated pairs but truncated one modality. For instance:
    • Truncated trajectory: $(\tau_{1},l_{1}+l_{2})$
    • Truncated instruction: $(\tau_{1}+\tau_{2},l_{1})$
    This assesses the model’s ability to detect partial mismatches in longer sequences.
A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
We list the figures of reward signals from learned VLMs for different types of trajectory-instruction pairs for each individual environment. All figures show that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards.
@@ A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma
@@
Table 7: Agent performance in Montezuma’s Revenge, evaluated across three rooms with the goal state being the exit through a designated door. Metrics measured are the Area Under the Curve (AUC) for total reward (where higher is better) and success rate (SR) of reaching the goal state. The baseline is denoted as B, with $\star$ marking statistical significance ($p<0.05$).
{tblr}
width = 0.95colspec = Q[473]Q[346]Q[106],row1 = p,hline1-2,8 = -,hline4,6 = -dashed Model & AUC SR
(1) PPO (Schulman et al., 2017) all failed 0%
(2) PPO+RND (B) (Burda et al., 2018) 0.550$\pm$0.066 100%
(3) PPO + Sim RM 1 (Perfect) 0.287$\pm$0.048 100%
(4) PPO+RND + Sim RM 1 (Perfect) 0.608$\pm$0.073 100%
(5) PPO+RND + Sim RM 2 (False Positive) 0.183$\pm$0.187 $\star$ 73.3%
(6) PPO+RND + Sim RM 3 (Temp. Insen.) 0.051$\pm$0.116 $\star$ 16.7%
  • Perfect auxiliary reward model have shown enhanced performance compared to weak RL baselines like PPO. While agents trained solely with PPO struggled to play Montezuma, incorporating Perfect auxiliary reward model into PPO (i.e., (3) in Table 7) did find the goal state, with the success rate increasing from 0% to 100%. However, we observed that the auxiliary reward model learns more slowly than the intrinsic reward model (see (2) in Table 7) when attempting to reach the goal state. This finding was initially surprising, but upon examining the agent movement heatmaps in Figure 15(a) and Figure 15(b), an explanation becomes clear: the PPO+RND model discovers shorter paths to the goal state. The heatmap reveals that the PPO+RND agent learns to directly jump to a rope, bypassing the use of a ladder and conveyor belt as suggested by the expert instructions. This observation raises a concern about instruction-following based reward signals; they might limit the agent’s exploration, confining it to states favored by the expert. If the quality of the expert’s knowledge about the task is not optimal, this can result in the agent failing to compete with an exploration-based reward model.
  • Again, we observed a remarkable synergy between the instruction-following-based reward model and the intrinsic reward model, as discussed in §7.1. As shown in (4) in Table 7, the PPO+RND + Sim RM 1 (Perfect) agent achieves the highest AUC score, demonstrating that it learns to reach the goal state faster than any other model.
  • The involvement of false positive rewards significantly slows down the learning process of the model, as evidenced by entries (5) and (6) in Table 7. Additionally, both (5) and (6) show lower AUC scores compared to (2), which is the vanilla PPO+RND model. If we consider the vanilla PPO+RND model to represent a scenario with full false negative instruction-following rewards (since no instruction-following reward model is implemented here), then comparing the scores of (2) with (5) and (6) further indicates that false positive rewards are more detrimental than false negative ones.
  • The RL + VLM reward approach completely failed to converge when the reward model disregarded temporal ordering, as evidenced by entry (6) in Table 7. The success rate had a significant drop to 16.7%, and even the RND intrinsic reward was unable to recover the success rate within the time budget. This illustrates the severe impact of false positive rewards due to temporal ordering insensitivity, an example of composition insensitivity discussed in §5; they trap the agent in a cycle of ineffective actions. However, note that if every sub-task is fulfilled, the agent will still receive the maximum total rewards, suggesting that, in theory, the policy could eventually converge. Yet, the practical implications of such false positives emphasize how they can drastically extend the learning period or lead to non-convergence in real-world scenarios where time or computational resources are limited.
(a) PPO+RND agent (same as + Sim RM (False Negative))
(b) PPO+RND + Sim RM 1 (Perfect)
(c) PPO+RND + Sim RM 2 (False Positive)
(d) PPO+RND + Sim RM 3 (Temporal Insensitivity)
We also provided movement heatmap of different models for further qualitative analysis. As shown in Figure 15(a), intrinsic reward model explored a shortcut towards the goal state by directly jump from the initial state to the rope, thereby learning to reach the goal state faster than other models.
Figures 15(b) and 15(c) do not exhibit a significant contrast, but they do reveal that the presence of false positive rewards leads the agent to more frequently visit dead ends, such as falling off cliffs. This behavior is consistent with the mechanics of the simulated false positive reward model, which defines a radius $\sigma$ around target states. Within this radius, even if an agent falls off a cliff, it might receive rewards because its position could be close enough to the intended intermediate rewarding state. Consequently, agents had a chance of mistakenly perceiving falling off as advantageous.
Figure 15(d) illustrates the consequences when temporal ordering restrictions are removed: the learning agent persistently pursues rewards associated with the final instruction, “walk left to the door.” However, simply executing this last instruction does not suffice for escaping the room because access to the door requires a key. This suggests that false positive rewards cause the agent to get trapped in local minima closer to $s_{0}$, indicating that deep RL can easily hack the reward system and effectively exploit shortcuts (Clark, 2016). A critical issue arises here: even with intrinsic rewards, the agent fails to escape these local optima, suggesting it has reached an equilibrium in the MDP where policy updates cease. This scenario is particularly relevant when the initial state distribution does not cover the entire state space. If the initial distribution does not have a positive probability for all states (i.e., $foralls\in\mathcal{S},d_{0}(s)>0$), it is not guaranteed that the RL algorithm will eventually reach the global optimum with maximum total rewards, as highlighted by (Agarwal et al., 2021).
A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
Using empirical quantile as threshold guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized while minimizing the average number of mistakes predicting false positives (Sadinle et al., 2019):
Algorithm 2 Calculate Empirical Quantile ($\hat{q}$)
Calibration set $\{\tau,l\}_{n}$, where $l$is the instruction sentence, $\tau$is the corresponding trajectory, and $n$is the number of samples; Significance level $\alpha$; VLM model reward model $v$Obtain the similarity-based score
[…]
2:$q_{\text{level}}\leftarrow\frac{\lceil(n-1)\times(1-\alpha)\rceil}{n}$LCommentCompute the empirical quantile
@@ A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
A.9 Implementation Details of the Experiments of BiMI Reward Function
We set confidence level for empirical quantile calculation to be $0.91-\alpha=$ 0.9. We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner, 2021). This constraint poses a significant challenge, particularly in sparse reward settings, as it demands that agents both explore efficiently and exploit their knowledge effectively within this limited budget.
To achieve the 1 million frame budget, we used the following configuration:
  • nproc: 8 (Number of processes used for parallel environments)
  • nstep: 512 (Length of the rollout stored in the buffer)
  • nepoch: 250 (Number of epochs to train the RL policy)
This configuration results in approximately 1 million steps: 250 epochs $\times$ 512 steps $\times$ 8 processes $=$ 1,024,000 frames.
In the case of Montezuma’s Revenge, we found that the 1 million frame limit used in Crafter was insufficient due to the game’s complexity and sparse reward structure. To address this, we extended the training budget to 8 million frames. It’s important to note that even with this increased frame count, agents were still unable to fully solve the task. As Zhang et al. (2021) pointed out, about 1 billion frames are required to truly master Montezuma’s Revenge. This vast difference in required training time (8 million vs 1 billion frames) underscores the exceptional difficulty of Montezuma’s Revenge as a sparse reward task.
The implementation details for the BiMI reward function are consistent with those outlined in the first stage of experiments.
A.10 Detailed Experiment Results of BiMI Reward Function
@@
A.10.1 Minigrid
ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid. The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge. Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals. The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge. This capability is crucial for deploying instruction-following agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
A.10.2 Crafter
We observed an intriguing pattern of results. The Bi component alone led to 14% and 3.2% improvement in performance over the original models. However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement. This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival. The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions. Furthermore, note that instruction-following RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter. This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture. Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments. The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
A.11 DEPRECATED Proof of the Reduction of Convergence Rate
WARNING: The following content is deprecated. This section contains the original convergence analysis for the ARR October Version, which has been identified as problematic by Reviewer nSDu. It is retained here solely for reference and backtracking purposes.
Formally, Abel et al. (2021) defined that:
Theorem A.7 (Abel et al. (2021)).
A reward function realizes a Range Set of Acceptable Policies (Range-SOAP) $\Pi_{G}$ when there exists an $\epsilon\geq 0$ such that every $\pi_{g}\in\Pi_{G}$ is $\epsilon$-optimal in start-state value, $V^{*}(s_{0})-V^{\pi_{g}}(s_{0})\leq\epsilon$, while all other policies are worse.
When the reward signal is sparse, the agent only receives a reward upon reaching some goal states, and the reward function does not provide any feedback during the intermediate steps. We argue that the sparse reward function realizes a Range-SOAP, leading to the categorization of policies into acceptable and unacceptable policies.
Proposition A.8.
Sparse reward function “realizes” Range-SOAP (i.e., Range Set of Acceptable Policies).
Justification:
A sparse reward function, where the agent only receives a reward upon completing the task, cannot prefer policies that lead to shorter task completion times. This is because either the agent completes the goal very quickly or slowly, they will receive nearly the same amount of cumulative rewards, and the reward function will not show a strong preference.
Since the sparse reward function does not induce a strict partial ordering on the policies, we say this reward function cannot realize a Partial Ordering (PO) task. Specifically, a PO on policies is a generalization of a Set of Acceptable Policies (SOAP) task. In a PO, the agent specifies a partial ordering on the policy space, where some policies are identified as “great”, some as “good”, and some as “bad” to strictly avoid, while remaining indifferent to the rest.
Therefore, the sparse reward function can realize a Set of Acceptable Policies (SOAP), where there is a set of policies that are all considered “good” or near-optimal, while all other policies are worse.
Furthermore, the sparse reward function will lead to a Range-SOAP, rather than an Equal-SOAP. Specifically, Equal-SOAP is a SOAP where all the acceptable policies are equally optimal in start-state value. This is because the good policies in the SOAP may differ slightly in their start-state values, as some may reach multiple goal states in the environment and thereby receiving different cumulative rewards. Therefore, the sparse reward function will realize a Range-SOAP, where there is a range of acceptable policies that are all near-optimal in start-state value.
This proposition forms the basis of our theoretical analysis in Section A.11.
Below is the convergence analysis when the total return can be classified by returns of trajectories from acceptable policies $P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$ versus returns of trajectories from unacceptable policies $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$, where $\mathcal{T}_{G}\sim\Pi_{G}$, $\mathcal{T}_{B}\sim\Pi_{B}$ and $\Pi_{B}=\Pi\setminus\Pi_{G}$.
Note on Derivation: The proof presented here can be viewed as a reformulation of the Q-value function’s gradient in the context of our specific problem setup. While this derivation may appear straightforward to readers well-versed in reinforcement learning, we include it to provide a clear mathematical foundation for our analysis of false positive rewards in the VLM-RL context. This formulation helps bridge the gap between standard RL theory and our specific problem domain on instruction-following RL with VLM-based rewards.
Specifically, the update rule of Actor-Critic algorithm is:
  • Critic:
    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (41)
    where
    $\delta=\mathbb{E}_{\pi_{\theta}}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]$ is the Monte Carlo (MC) estimation error, $Q_{\phi}(s,a)$ is the $Q$-value function which measures the expected discounted cumulative reward given the state $s$ and action $a$, and $\pi_{\theta}$ is the policy.
    $G^{\pi_{\theta}}$ is the rollout cumulative rewards from the trajectory $\tau^{\pi_{\theta}}$ generated from $\pi_{\theta}$.
  • Actor:
    $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ (42)
We need to make the following assumptions to simplify the theoretical analysis:
  1. 1.
    We use $\mathcal{T}_{G}$ to represent the set of acceptable trajectories and use $\mathcal{T}_{B}$ as a non-overlapping set of “not acceptable” trajectories, we can say:
    $\mathcal{T}_{G}\cap\mathcal{T}_{B}=\emptyset$ (43)
    where $\mathcal{T}$ is the trajectory distribution space of the agent policy $\pi_{\theta}$.
  2. 2.
    When we define $\mathcal{T}_{B}$ and $\mathcal{T}_{G}$ within the universe $\mathcal{T}$, we must account for the possibility of a remaining set to ensure a complete partition of $\mathcal{T}$. The remaining set $\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\}$ can be interpreted as a trajectories that partially fulfill the instruction. Given our focus on the influence of false positive rewards within challenging sequential decision-making tasks, we assume that partially correct trajectories are also unacceptable. This stance is justified by the fact that partially correct trajectories can be viewed as false positives: they appear to be plausible behaviors but ultimately fail to fulfill the instructions. Rewarding such trajectories could reinforce incomplete or potentially dangerous behaviors. The errors discussed in Section 5 can be seen as instances of partially correct trajectories that are ultimately unacceptable. For example, incompletely following a safety-critical task such as “turn off the machinery before performing maintenance” could pose significant risks, and rewarding such behavior would be undesirable. Consequently, for the purposes of our analysis, we can assume:
    $\mathcal{T}_{G}\cup\mathcal{T}_{B}=\mathcal{T}$ (44)
  3. 3.
    Assume the policy class parameterized by $\theta$ should be expressive enough to capture optimal or near-optimal policies, and the policy is initialized randomly from uniform distribution, i.e., $\pi_{\theta_{0}}\sim\mathcal{U}(\Pi)$. Meanwhile, the Q-value is initialized as $Q_{\phi_{0}}(s,a)=0,\forall_{s\in S}\forall_{a\in A}$.
  4. 4.
    We assume that $|\mathcal{T}_{G}|$ and $|\mathcal{T}_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)-Q_{\phi}(s,a)]\|^{2}$, the updated $Q$-value will approach as follows:
$Q_{\phi}(s,a)\rightarrow\mathbb{E}[G(\tau)]$ (45)
$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]+P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$
$+P(\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})]$ (46)
[…]
Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$, we have the following:
$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$ (49)
$=$ $\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\pi_{\theta}(a\mid s)\big{)}$
$+\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\pi_{\theta}(a\mid s)\big{)}$ (50)
$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)\quad\text{$\theta_{\text{old}}$ terms are constants w.r.t. $\theta$}$ (51)
$=$ $\big{(}1-P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\big{)}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ (52)
$=$ $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
[…]
Note the highlighted term in green ($P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})$) is the direction of the gradient ascent on parameter $\theta$.
Justification of the assumptions.
  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al. (2021)).
  • In reference to the realizability condition implied by Assumption 3, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al., 1989).
Observation. Since the goal of the learning agent is to maximize $P(\tau\in\mathcal{T}_{G}\mid\theta)$ (i.e., to converge the agent policy to the distribution of acceptable trajectories), we can see that the second term provides the target direction with rate $(\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]-\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}])$. Therefore, the ascent rate will decrease when the cumulative rewards from unacceptable trajectories (i.e., the false positive rewards) gets higher. In addition, the first term $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ can be regarded as the deviation term of target direction. This formalization follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.

### v5

@@ H.3 Hyperparameters for VLM Reward Model + RL Agents
   - chans [64, 128, 128]
   - outsize 256
   - nblock 2
   - post_pool_groups 1
   - init_norm_kwargs
   - batch_norm false
   - group_norm_groups 1
@@
   - layer_norm true
Table 6: Crafter and Minigrid RL Training Parameters
@@
   - ppo_nepoch 3
   - ppo_nbatch 8
   - clip_param 0.2
   - vf_loss_coef 0.5
   - ent_coef 0.01
   - lr 3.0e-4
   - max_grad_norm 0.5
   - aux_freq 8
   - aux_nepoch 6
   - pi_dist_coef 1.0
   - vf_dist_coef 1.0
Table 7: Montezuma’s Revenge RL Training Parameters
@@
   - update_proportion 0.25
   - ppo_nepoch 3
   - ppo_batch_size 256
   - clip_param 0.1
   - vf_loss_coef 0.5
   - ent_coef 0.001
   - lr 1.0e-4
H.4 Evaluation Metrics
As detailed in Appendix A.1, we adopted the score metric from the Crafter benchmark (Hafner 2022) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments. Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress.
Appendix I Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
In this work, we have shown that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards. We list the figures of reward signals from learned VLMs for different types of trajectory-instruction pairs for each individual environment — the Montezuma, Minigrid, and Crafter environments. All figures show that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards.
@@ Appendix I Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
Appendix J Extra Experiment on Simulated Oracle Reward Model to demonstrate H3 and H4
@@
Table 8: Agent performance in Montezuma’s Revenge, evaluated across three rooms with the goal state being the exit through a designated door. Metrics measured are the Area Under the Curve (AUC) for total reward (where higher is better) and success rate (SR) of reaching the goal state. The baseline is annotated using a $\star$ symbol. The results are averaged over 5 runs, with standard deviations reported.
Model AUC SR
(1) PPO (Schulman et al. 2017) all failed 0%
@@
[…]
  1. 1.
    Perfect auxiliary reward model have shown enhanced performance compared to weak RL baselines like PPO. While agents trained solely with PPO struggled to play Montezuma, incorporating Perfect auxiliary reward model into PPO (i.e., entry (3) in Table 8) did find the goal state, with the success rate increasing from 0% to 100%. However, we observed that the auxiliary reward model learns more slowly than the intrinsic reward model. This is evidenced by its lower AUC score (0.287 for entry (3) vs. 0.550 for entry (2) in Table 8), where AUC reflects how quickly the agent learn to reach the goal state. This finding was initially surprising, but upon examining the agent movement heatmaps in Figure 15(a) and Figure 15(b), an explanation becomes clear: the PPO+RND model discovers shorter paths to the goal state. The heatmap reveals that the PPO+RND agent learns to directly jump to a rope, bypassing the use of a ladder and conveyor belt as suggested by the expert instructions. This observation highlights a nuance in VLM-based reward signals, potentially overlooking more efficient routes when expert instructions do not serve as landmarks (a well-known concept in heuristic search in the field of classical planning).
  2. 2.
    We observed a remarkable synergy between the instruction-based reward model and the intrinsic reward model. As shown in entry (4) in Table 8, the PPO+RND + Sim RM 1 (Perfect) agent achieves the highest AUC score, demonstrating that it learns to reach the goal state faster than any other model.
  3. 3.
    False positive rewards significantly impeded learning, as shown by entries (5) and (6) in Table 8. Heatmaps (Figure 15(b), Figure 15(c)) reveal agents frequently visiting dead ends (e.g., falling off cliffs) under these conditions, suggesting that the agent is trapped in local minima. To assess their impact, we compare these to PPO+RND (entry (2)), a baseline lacking VLM-based instruction rewards. Since (2) does not even have a VLM-based reward model, we treat it as a baseline with full false negative instruction-based rewards. The AUC scores of (5) and (6) are lower than (2), indicating that false positives are more detrimental than false negatives, supporting our hypothesis H4, as detailed shortly.
  4. 4.
    When the VLM reward model ignored temporal ordering (entry (6)), convergence failed, with the success rate dropping to 16.7% and RND unable to recover within the time limit. Heatmap (Figure 15(d)) illustrates the agent fixating on the final instruction (“walk left to the door”) without securing the key, trapping it in local minima near $s_{0}$ due to false positives — a form of composition insensitivity (Section 5). Moreover, given that Montezuma’s Revenge environment has a fixed initial state distribution, (i.e., the agent always starts at a fixed $s_{0}$), it is not guaranteed that the RL algorithm will eventually reach the global optimum with maximum total rewards, as highlighted by Agarwal et al. (2019).
(a) PPO+RND (same as + RM (False Negative))
(b) PPO+RND + Sim RM 1 (Perfect Simulation)
(c) PPO+RND + Sim RM 2 (False Positive)
(d) PPO+RND + Sim RM 3 (Temporal Insensitivity)
Figure 15 provides initial evidence of the differential impact of false negatives versus false positives on training outcomes, as posited by H4, suggesting that false positives — particularly those tied to temporal insensitivity — more severely degrade final scores.
Appendix K Algorithm for Empirical Quantile Calculation
The empirical quantile $\hat{q}$ is calculated using the conformal prediction method, shown as follows:
Algorithm 2 Calculate Empirical Quantile ($\hat{q}$) in VLM RM
1: Calibration set $\{\tau,l\}_{n}$, where $l$ is the instruction sentence, $\tau$ is the corresponding trajectory, and $n$ is the number of samples; User-defined error rate $\alpha$; VLM model reward model $v$
[…]
3: $r\}_{n}\leftarrow\{v(\tau,l)\}_{n}$
[…]
5: $q_{\text{level}}\leftarrow\frac{\lceil(n-1)\times(1-\alpha)\rceil}{n}$
[…]
7: $\hat{q}\leftarrow\texttt{np.quantile($\{r\}_{n}$, $q_{\text{level}}$, method=`lower')}$
8: return $\hat{q}$
Appendix L Ablation Study Lineplots for BiMI reward
@@
Appendix M Encoder-based Vision-Language Models Explanation
Similar to LLMs, VLMs can be broadly classified into two categories: encoder-based models (or dual-encoder models) and generative-based models (often referred to as Vision LLMs).
Encoder-based models focus on mapping vision and language data into a shared semantic space where embeddings of corresponding visual and textual content are positioned close to each other. CLIP (Radford et al. 2021) and SigLIP (Tschannen et al. 2025) exemplify this approach. They are often referred to as VLM dual-encoders as such VLMs consists of both vision encoder and text encoder which can be operated separately to encode vision or language data. Dual-encoder models are particularly well-suited for cross-modal retrieval tasks, such as finding the most relevant image for a given text query or vice versa. This is achieved by computing similarity scores (e.g., cosine similarity) or other distance metrics between the embeddings of the query and the images in the dataset. Dual-encoders are therefore being explored to serve as learned-based reward models for embodied RL agents (Rocamonde et al. 2024; Du et al. 2023), which we discuss in Appendix B.
In contrast, generative-based models such as LLaVA (Liu et al. 2023) and Sa2VA (Yuan et al. 2025) are typically built by connecting a vision encoder (e.g., CLIP-based vision encoder) to an LLM via a projection layer (e.g., a Multilayer Perceptron (MLP)). The backbone LLM takes both vision feature vectors and text tokens as input and generates textual outputs, enabling these models to perform tasks such as Vision Question Answering (VQA). Although both categories are commonly classified as VLMs, they serve fundamentally different purposes and exhibit distinct behaviors.
Appendix N Sparse Reward and Random Walk
Below, we will discuss the characteristics of RL training under sparse rewards
Proposition N.1.
With sparse rewards, the gradient landscape is nearly flat, making gradient-ascent updates indistinguishable from a random walk in parameter space.
While this may seem obvious, no existing textbook has yet explained this concept clearly. Here, we aim to formalize and justify this proposition, as it will be crucial for analyzing the performance of VLM-based reward models in this work.
Recall the gradient of the policy gradient update in online Actor Critic methods: $\theta\leftarrow\theta+\alpha\mathbb{E}_{\pi_{\theta}}\left[Q_{\phi}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\right]$. Using chain rule, the gradient of the policy gradient update can be decomposed as:
$\mathbb{E}_{\pi_{\theta}}\left[\frac{Q_{\phi}(s,a)}{\pi_{\theta}(a|s)}\nabla_{\theta}\pi_{\theta}(a|s)\right]$
In sparse reward settings, the Q-function $Q^{\pi_{\theta}}(s,a)$, often approximated by a neural network in actor-critic methods, exhibits significant randomness. Initially, the critic’s weights are randomly initialized (e.g., via Xavier initialization (Glorot and Bengio 2010)), producing arbitrary Q-values (Mnih et al. 2015). With most transitions yielding $r=0$, the TD error $\delta=\gamma Q(s^{\prime},a^{\prime})-Q(s,a)$ is typically non-zero due to the stochastic variation between $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ resulting from random initialization. This leads to random updates to the Q-network, as the noisy TD error lacks a consistent direction, further increasing the stochasticity of Q-value estimates.
Such stochasticity propagates to the policy gradient update, where it even amplifies this effect: At initialization, $\pi_{\theta}(a|s)\approx 1/K$ for a discrete action space with $K$ actions, as random weights yield a near-uniform softmax distribution. The gradient term $\frac{\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ scales $Q_{\phi}(s,a)$ by the inverse of the policy probability. Therefore, the update step $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ has a high variance due to noisy Q-value estimates and non-trivial gradient magnitudes from the scaling effect of $\pi_{\theta}(a|s)$. Consequently, the update exhibits $\text{Var}(\hat{\nabla}_{\theta}J(\theta))>0$ and $\mathbb{E}[\hat{\nabla}_{\theta}J(\theta)]\approx 0$.
This behavior renders policy updates indistinguishable from a random walk in parameter space. The parameters $\theta$ drift stochastically, as each step’s direction is dominated by noise rather than progress toward a better policy. This random walk behavior is a key reason why training under sparse rewards is challenging.